OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = –x6 + x5– x4 + 4x3 – 12x2 + 12 ?

OpenStudy (lgbasallote):

use descarte's rule of signs f(x) = (-) (+)( -) (+) (-) (+) 5 change in signs so 5 possible positive roots f(-x) = (-) (-) (-) (-) (-) (+) 1 change in sign so 1 possible negative root 5 positive roots + 1 negative root = 6 roots since 6 is the degree of this equation there are only 6 possible roots and since they're already occupied there are no complex roots does that help?

OpenStudy (anonymous):

All the numbers at x are sqaured

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

–x^6 + x^5– x^4 + 4x^3 – 12x^2 + 12 ?

OpenStudy (lgbasallote):

yeah...exactly...

OpenStudy (lgbasallote):

...you have no idea what the descartes' rule of signs is...?

OpenStudy (anonymous):

I'm trying to understand it.

OpenStudy (lgbasallote):

as the name suggests...you're going to look at the signs...then you search for change in signs

OpenStudy (lgbasallote):

for every change in sign in f(x) counts for 1 possible positive root every change in sign in f(-x) counts for 1 possible negative root

OpenStudy (anonymous):

Okay

OpenStudy (lgbasallote):

just to be clear... f(-x) = -(-x)^6 + (-x)^5 - (-x)^4 + 4(-x)^3 - 12(-x)^2+12 f(-x) = -(x^6) + (-x^5) -(x^4) + 4(-x^3) -12(x^2) + 12 f(-x) = -x^6 - x^5 - x^4 - 4x^3 -12x^2 + 12

OpenStudy (lgbasallote):

i'll let you understand it now :) tag me if you have questions

OpenStudy (anonymous):

@lgbasallote bravo